# Customizing Text-to-Image Models with a Single Image Pair

### v1

@@ 1 Introduction
Figure 1: Given a single image pair, we present Pair Customization, a method for customizing a pre-trained text-to-image model and learning a new style from the image pair’s stylistic difference. Our method can apply the learned stylistic difference to new input images while preserving the input structure. Compared to Dreambooth LoRA[33, 73], a standard customization method that solely use style images, our method effectively disentangles style and content, resulting in better structure, color preservation, and style application. Style image credit: Jack Parkhouse.
@@ 2 Related Works
Customizing generative models.   Model customization, or personalization, aims to adapt an existing generative model with additional data, with the goal of generating outputs tailored to specific user preferences. Earlier efforts mainly focus on customizing pre-trained GANs [25, 40, 41] for smaller datasets [39, 98, 57], incorporating user edits [90, 91, 6], or aligning with text prompts [21, 58]. Recently, the focus has pivoted towards adapting large-scale text-to-image models to generate user-provided concepts, typically presented as one or a few images. Simply fine-tuning on the concept leads to overfitting. To mitigate this and enable variations via free text, several works explored different regularizations, including prior preservation [71, 43], human alignment [82], as well as parameter update restriction, where we only update text tokens [18, 15, 1, 89], attention layers [43, 19, 27], low-rank weights [33, 86, 73], or clusters of neurons [50]. More recent methods focus on encoder-based approaches for faster personalization [2, 20, 94, 44, 13, 88, 72, 53, 14, 80, 95]. Several works further focus on multiple concepts [43, 64, 26, 79, 3] instead of only a single concept. Our method takes inspiration from these techniques; however, we aim to address an inherently different task. Instead of learning concepts from an image collection, we customize the model to learn stylistic differences from an image pair.
Style and content separation.   Various past works have explored learning a style while separating it from content [85, 23, 45, 34, 12]. Our work is inspired by the seminal work Image Analogy [29], a computational paradigm that takes an image pair and applies the same translation to unseen images. Common image analogy methods include patch-wise similarity matching [29, 36, 47] and data-driven approaches [68, 87, 99, 60, 5, 92]. Different from these, we aim to exploit the text-guided generation capabilities of large-scale models so that we can directly use the style concept with unseen context. Recently, StyleDrop [82] has been proposed to learn a custom style for masked generative transformer models. Concurrent with our work, Hertz et al.[28] introduced a method for generating images with style consistency, offering the option of using a style reference image. In contrast, we exploit an image pair to better discern the stylistic difference. Unlike StyleDrop, we do not rely on human feedback in the process.
@@ 3.1 Preliminary: Model Customization
Diffusion models.   Diffusion models[30, 84, 81], map Gaussian noise to the image distribution through iterative denoising. Denoising is learned by reversing the forward diffusion process ${\mathbf{x}}_{0},\dots,{\mathbf{x}}_{T}$, where image ${\mathbf{x}}_{0}$ is slowly diffused to random noise ${\mathbf{x}}_{T}$ over $T$ timesteps, defined by ${\mathbf{x}}_{t}=\sqrt{\widebar{\alpha}_{t}}{\mathbf{x}}_{0}+\sqrt{1-\widebar{\alpha}_{t}}\epsilon$ for timestep $t\in[0,T]$. Noise $\epsilon\sim\mathcal{N}(0,I)$ is randomly sampled, and $\widebar{\alpha}_{t}$ controls the noise strength. The training objective of diffusion models is to denoise any intermediate noisy image ${\mathbf{x}}_{t}$ via noise prediction:
@@ 3.2 Style Extraction from an image pair
We aim to customize a pre-trained model with an artistic style in order to stylize the original model outputs while preserving their content, as shown in Figure 2 (right). To achieve this, we introduce style LoRA weight $\theta_{\text{style}}=\theta_{0}+\Delta\theta_{\text{style}}$. While a pre-trained model generates content from a noise seed and text $c$, style LoRA’s goal is to generate a stylized counterpart of original content from the same noise seed and a style-specific text prompt ${\mathbf{c}}_{\text{style}}$, where ${\mathbf{c}}_{\text{style}}$ is original text $c$ appended by suffix “in <desc> style”. Here, <desc> is a placeholder for some worded description of the style (e.g., “digital art”), and style LoRA $\theta_{\text{style}}$ associates <desc> to the desired style.
@@
Disentangling style and content.   We leverage the fact that the style image share the same layout and structure as the content image. Our key idea is to learn a separate content LoRA $\theta_{\text{content}}=\theta_{0}+\Delta\theta_{\text{content}}$ to reconstruct the content image. By explicitly modeling the content, we can train the style LoRA to “extract” the stylistic differences between the style and content image. We apply both style and content LoRA to reconstruct the style image, i.e., $\theta_{\text{combined}}=\theta_{0}+\Delta\theta_{\text{content}}+\Delta\theta_{\text{style}}$. This approach prevents leaking the content image to style LoRA, resulting in a better stylization model.
During training, we feed the content LoRA $\theta_{\text{content}}$ with a content-specific text ${\mathbf{c}}_{\text{content}}$, which contains a random rare token V*, and feed the combined model $\theta_{\text{combined}}$ with ${\mathbf{c}}_{\text{style}}$, where ${\mathbf{c}}_{\text{style}}$ is “$\{{\mathbf{c}}_{\text{content}}\}$ in <desc> style”. Figure 2 (Left) summarizes our training process.
@@
where $\epsilon_{\theta_{\text{combined}}}$ is the denoiser with both LoRAs applied as in Equation 5, ${\mathbf{c}}_{\text{style}}$ is “$\{{\mathbf{c}}_{\text{content}}\}$ in <desc> style”, and <desc> is a worded description of the style (e.g., “digital art”). Finally, we jointly optimize the LoRAs with the two losses:
@@
Figure 3: (Left) Orthogonal adaptation. Enforcing row-space orthogonality between style and content LoRA improves image quality, where the images capture the style better and have fewer visual artifacts. (Right) Style guidance. We compare style control capabilities between our style guidance and standard LoRA weight scaling [74]. Blue and green stand for the LoRA weight scale and style guidance scale, respectively. Style guidance better preserves content when the style is applied. More details of style guidance formulation are in Section 3.3.
@@
$W_{\text{content}}=W_{0}+B_{\text{content}}A_{\text{content}};\;W_{\text{style}}=W_{0}+B_{\text{style}}A_{\text{style}}.$ (8)
@@ 3.3 Style Guidance
$\hat{\epsilon}_{\theta_{0},\theta_{\text{style}}}({\mathbf{x}}_{t},{\mathbf{c}},{\mathbf{c}}_{\text{style}})=$ $\epsilon_{\theta_{0}}({\mathbf{x}}_{t},\varnothing)$ (10)
@@
where style guidance is the difference in noise prediction between style LoRA and the pre-trained model. Style guidance strength is controlled by $\lambda_{\text{style}}$, and setting $\lambda_{\text{style}}=0$ is equivalent to generating original content. In Figure 3, we compare our style guidance against scaling LoRA weights (Equation 3), and we find that our method better preserves the layout. More details and a derivation of our style guidance are in Appendix 0.B.
Previous works [49, 8] have also considered applying multiple guidance terms with diffusion models. A major difference in our approach is that we obtain additional guidance from a customized model and apply it to the original model. Styledrop [82] considers a similar formulation with two guidance terms but for transformer-based generative models.
Blending multiple learned styles.   With a collection of models customized by our method, we can blend the learned styles as follows. Specifically, given some set of styles $\mathcal{S}$ and strengths $\lambda_{\text{style}_{0}},\dots,\lambda_{\text{style}_{n}}$, we can blend the style guidance from each model, and our new inference path is represented by
$\hat{\epsilon}_{\theta_{0},\theta_{\text{style}}}({\mathbf{x}}_{t},{\mathbf{c}},{\mathbf{c}}_{\text{style}})=$ $\epsilon_{\theta_{0}}({\mathbf{x}}_{t},\varnothing)$ (11)
@@ 4.1 Dataset
Figure 4: Quantitative comparison with baselines on learned style. Given a fixed inference path, our method’s pareto dominates baselines for image generation both on the same category as training (left) and when evaluated on categories different from training, e.g., trained on human portraits but tested on dog images (right). Secondly, our proposed style guidance outperforms standard LoRA weight scale guidance for our training method (blue vs. orange), DB LoRA (green vs. pink), and Sliders (brown vs. purple). In the Appendix’s Figure 9, we further evaluate the diversity of generated images. We show that baselines often lose diversity, while our method leads to diverse generations while still achieving lower perceptual distance to the ground truth style. Increased marker size corresponds to an increase in style guidance scale.
Datasets.   To enable large-scale quantitative evaluation, we construct a diverse set of paired style and content images as follows. First, we generate 40 content images for each class: headshots, animals, and landscapes. When generating images in the headshot class, we generate 20 images with the prompt “A professional headshot of a man” and 20 images with the prompt “A professional headshot of a woman”. Similarly, we split the animal class into photos of dogs and cats. To curate synthetic pairs, we then apply image editing or image-to-image translation methods to all the content images to obtain the stylized version. For each unique prompt, we choose a single paired instance as training data and hold out the other pairs with the same prompt as a test set (Same Category). For each prompt, we also choose 5 pairs from each of the other prompts as a secondary test set (Different Category). We show all our synthetic training image pairs in Appendix 0.C. By leveraging synthetic pairs for evaluation, we can train on a single synthetic pair and test our results against held out synthetic style images. Secondly, we qualitatively compare against single artist pairs in Figure 6. Next, we describe the specific methods to create the paired dataset.
@@ 4.2 Baselines and Evaluation Metrics
Baselines.   We compare our method against – (1) DreamBooth LoRA [33, 74] (DB LoRA), (2) Concept Sliders [22] (3) IP-adapters [95], (4) IP-adapters w/ T2I, and (5) StyleDrop [82]. DB LoRA uses only the style image and fine-tunes low-rank adapters in all the linear layers in the attention blocks of the diffusion model. We evaluate different amounts of style applications for DB LoRA using the standard LoRA scale [74] and our style guidance. Concept sliders presents a paired image model customization method that trains a single low-rank adapter jointly on both images, with different reconstruction losses for the style and content images. We also evaluate using both the standard LoRA scale and our style guidance. IP-adapters is an encoder-based method that does not require training for every style and takes a style image as an extra condition separate from the text prompt. Increasing or decreasing the guidance from the input style image is possible by scaling the weight of the image conditioning. We consider the SDXL [65] implementation of this method. For the IP-Adapter, we also compare against the stronger baseline of providing extra conditioning of an edge map of the content image through T2I Adapters [56] to preserve the content image structure. The recently proposed Styledrop [82] technique for learning new styles is based on MUSE [11], and uses human feedback in its method. Since MUSE is not publicly available, we follow Style-Aligned Image Generation’s [28] setup, and implement a version of StyleDrop on SDXL. Specifically, we train low-rank linear layers following each Feed-Forward layer in the attention blocks of SDXL. For a fair comparison, we train Styledrop without human feedback.
@@
Evaluation metrics.   When evaluating the performance of each method, we consider two quantitative metrics: perceptual distance to ground truth style images and structure preservation from the original image. A better customization method will have a low perceptual distance to the ground truth style images while still preserving content of the original image before adding style. We measure these using – (1) Distance to GT Styled: given holdout ground truth style images, we measure the perceptual distance between our styled outputs and the ground truth style images using DreamSim [17], a recent method for measuring the perceptual distance between images. DreamSim image embeddings are comprised of an ensemble of image embedding models, including CLIP [66] and DINO [10], which are then fine-tuned so the final embeddings respect human perception. We measure DreamSim distance as (1 - cosine similarity) between DreamSim embeddings, where a lower value implies that the images are perceptually more similar. (2) Distance to Content Image: to measure content preservation after style application, we measure the perceptual distance of our generated style image to the original content image with no style guidance. We again use DreamSim, this time comparing styled and content images. Note here that a perceptual distance of zero to the content image is undesirable, as this would require no style to be applied. However, a better-performing method should obtain a better tradeoff between the two distances. (3) We also perform a human preference study of our method against baselines.
@@
Figure 6: Result of our method compared to the strongest baselines. When only training with the style image as in DB LoRA, the image structure is not preserved and overfitting occurs. While Concept Slider’s training scheme [22] uses both style and content images, it still exhibits overfitting and loss of structure in many cases. Our method preserves the structure of the input mage while faithfully applying the desired style. We use style guidance strength 3 and classifier guidance strength 5. Style image credits: Jack Parkhouse (Third row) and Aaron Hertzmann (Fourth row)

### v2

@@ Abstract
Figure 1: Given a single image pair, we present Pair Customization, a method for customizing a pre-trained text-to-image model and learning a new style from the image pair’s stylistic difference. Our method can apply the learned stylistic difference to new input images while preserving the input structure. Compared to Dreambooth LoRA[36, 77], a standard customization method that solely uses style images, our method effectively disentangles style and content, resulting in better structure, color preservation, and style application. Style image credit: Jack Parkhouse.
@@ 2 Related Works
Customizing generative models.   Model customization, or personalization, aims to adapt an existing generative model with additional data, with the goal of generating outputs tailored to specific user preferences. Earlier efforts mainly focus on customizing pre-trained GANs [27, 43, 44] for smaller datasets [42, 104, 61], incorporating user edits [95, 96, 6], or aligning with text prompts [22, 62]. Recently, the focus has pivoted towards adapting large-scale text-to-image models to generate user-provided concepts, typically presented as one or a few images. Simply fine-tuning on the concept leads to overfitting. To mitigate this and enable variations via free text, several works explored different regularizations, including prior preservation [75, 46], human alignment [86], patch-based learning [103], as well as parameter update restriction, where we only update text tokens [19, 15, 1, 94], attention layers [46, 20, 30], low-rank weights [36, 90, 77], or clusters of neurons [53]. More recent methods focus on encoder-based approaches for faster personalization [2, 21, 99, 47, 13, 92, 76, 56, 14, 84, 100]. Instead of learning a single concept, several works further focus on learning multiple concepts [46, 67, 28, 83, 3]. Other methods [72, 58] propose customizing text-to-video models to learn motion, while Guo et al. [29] propose animating customized text-to-image models models by incorporating motion Low-Rank Adapter [36] modules. Our method takes inspiration from these techniques; however, we aim to address an inherently different task. Instead of learning concepts from an image collection, we customize the model to learn stylistic differences from an image pair.
Style and content separation.   Various past works have explored learning a style while separating it from content [89, 25, 49, 37, 12]. Our work is inspired by the seminal work Image Analogy [32], a computational paradigm that takes an image pair and applies the same translation to unseen images. Common image analogy methods include patch-wise similarity matching [32, 39, 50] and data-driven approaches [71, 91, 105, 63, 5, 97]. Different from these, we aim to exploit the text-guided generation capabilities of large-scale models so that we can directly use the style concept with unseen context. Recently, StyleDrop [86] has been proposed to learn a custom style for masked generative transformer models. Unlike StyleDrop, we do not rely on human feedback in the process. Concurrent with our work, Hertz et al.[31] introduced a method for generating images with style consistency, offering the option of using a style reference image. In contrast, we exploit an image pair to better discern the stylistic difference.
@@ 3.1 Preliminary: Model Customization
Diffusion models.   Diffusion models[33, 88, 85], map Gaussian noise to the image distribution through iterative denoising. Denoising is learned by reversing the forward diffusion process ${\mathbf{x}}_{0},\dots,{\mathbf{x}}_{T}$, where image ${\mathbf{x}}_{0}$ is slowly diffused to random noise ${\mathbf{x}}_{T}$ over $T$ timesteps, defined by ${\mathbf{x}}_{t}=\sqrt{\bar{\alpha}_{t}}{\mathbf{x}}_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon$ for timestep $t\in[0,T]$. Noise $\epsilon\sim\mathcal{N}(0,I)$ is randomly sampled, and $\bar{\alpha}_{t}$ controls the noise strength. The training objective of diffusion models is to denoise any intermediate noisy image ${\mathbf{x}}_{t}$ via noise prediction:
@@ 3.2 Style Extraction from an image pair
We aim to customize a pre-trained model with an artistic style in order to stylize the original model outputs while preserving their content, as shown in Figure 2 (right). To achieve this, we introduce style LoRA weight $\theta_{\text{style}}=\theta_{0}+\Delta\theta_{\text{style}}$. While a pre-trained model generates content from a noise seed and text $c$, style LoRA’s goal is to generate a stylized counterpart of original content from the same noise seed and a style-specific text prompt ${\mathbf{c}}_{\text{style}}$, where ${\mathbf{c}}_{\text{style}}$ is original text $c$ appended by suffix “in ¡desc¿ style”. Here, ¡desc¿ is a placeholder for some worded description of the style (e.g., “digital art”), and style LoRA $\theta_{\text{style}}$ associates ¡desc¿ to the desired style.
@@
Disentangling style and content.   We leverage the fact that the style image shares the same layout and structure as the content image. Our key idea is to learn a separate content LoRA $\theta_{\text{content}}=\theta_{0}+\Delta\theta_{\text{content}}$ to reconstruct the content image. By explicitly modeling the content, we can train the style LoRA to “extract” the stylistic differences between the style and content image. We apply both style and content LoRA to reconstruct the style image, i.e., $\theta_{\text{combined}}=\theta_{0}+\Delta\theta_{\text{content}}+\Delta\theta_{\text{style}}$. This approach prevents leaking the content image to style LoRA, resulting in a better stylization model.
During training, we feed the content LoRA $\theta_{\text{content}}$ with a content-specific text ${\mathbf{c}}_{\text{content}}$, which contains a random rare token V*, and feed the combined model $\theta_{\text{combined}}$ with ${\mathbf{c}}_{\text{style}}$, where ${\mathbf{c}}_{\text{style}}$ is “$\{{\mathbf{c}}_{\text{content}}\}$ in ¡desc¿ style”. Figure 2 (Left) summarizes our training process.
@@
where $\epsilon_{\theta_{\text{combined}}}$ is the denoiser with both LoRAs applied as in Equation 5, ${\mathbf{c}}_{\text{style}}$ is “$\{{\mathbf{c}}_{\text{content}}\}$ in ¡desc¿ style”, and ¡desc¿ is a worded description of the style (e.g., “digital art”). Finally, we jointly optimize the LoRAs with the two losses:
@@
Figure 3: Style guidance. We compare our style guidance and standard LoRA weight scaling [78]. Style guidance better preserves content when the style is applied. Blue and green stand for the LoRA weight scale and style guidance scale, respectively. More details of style guidance formulation are in Section 3.3.
Figure 4: Orthogonal adaptation. Enforcing row-space orthogonality between style and content LoRA improves image quality, where the images capture the style better and have fewer visual artifacts.
@@
$W_{\text{content}}=$ $W_{0}+B_{\text{content}}A_{\text{content}}$ (8)
$\;W_{\text{style}}=$ $W_{0}+B_{\text{style}}A_{\text{style}}.$
@@ 3.3 Style Guidance
$\hat{\epsilon}_{\theta_{0},\theta_{\text{style}}}$ $({\mathbf{x}}_{t},{\mathbf{c}},{\mathbf{c}}_{\text{style}})$ (10)
$=$ $\epsilon_{\theta_{0}}({\mathbf{x}}_{t},\varnothing)$
@@
where style guidance is the difference in noise prediction between style LoRA and the pre-trained model. Style guidance strength is controlled by $\lambda_{\text{style}}$, and setting $\lambda_{\text{style}}=0$ is equivalent to generating original content. In Figure 3, we compare our style guidance against scaling LoRA weights (Equation 3), and we find that our method better preserves the layout. More details and a derivation of our style guidance are in Appendix B.
Previous works have also used multiple guidance terms with diffusion models, including guidance from multiple text prompts using the same model [52] and additional image conditions [8]. Unlike these, we obtain additional guidance from a customized model and apply it to the original model. StyleDrop [86] considers a similar formulation with two guidance terms but for masked generative transformers. SINE [103] uses a customized content model to apply text-based image editing to a single image, like adding snow. However, we use a customized style model to generate any image with the desired style.
Blending multiple learned styles.   With a collection of models customized by our method, we can blend the learned styles as follows. Given some set of styles $\mathcal{S}$ and strengths $\lambda_{\text{style}_{0}},\dots,\lambda_{\text{style}_{n}}$, we blend the style guidance from each model, and our new inference path is represented by
$\hat{\epsilon}_{\theta_{0},\theta_{\text{style}}}$ $({\mathbf{x}}_{t},{\mathbf{c}},{\mathbf{c}}_{\text{style}})$ (11)
$=$ $\epsilon_{\theta_{0}}({\mathbf{x}}_{t},\varnothing)$
@@
Figure 5: Result of our method compared to the strongest baselines. When only training with the style image as in DB LoRA, the image structure is not preserved, and overfitting occurs. While Concept Slider’s training scheme [23] uses both style and content images, it still exhibits overfitting and loss of structure in many cases. Our method preserves the structure of the input image while faithfully applying the desired style. We use style guidance strength $3$ and classifier guidance strength $5$. Style image credits: Jack Parkhouse (First row) and Aaron Hertzmann (Second row)
Figure 6: Result of our method compared to the methods without finetuning (zoom in for best viewing). For the baseline methods, we also add the edge map from the pretrained output as an extra condition ($3^{\text{rd}}$ and $5^{\text{th}}$ column). Without this edge map, other methods tend to lose the structure of the pretrained output. In some cases, however, an additional edge map can overly constrain the output of a model, like in the second image pair example. Our method preserves the structure of the Stable Diffusion image while faithfully applying the desired style. We use style guidance strength $3$ and classifier guidance strength $5$ for our method and set the IP-adapter scale and style-alignment scale to $0.5$.
@@ 4.1 Dataset
Figure 7: Quantitative comparison with baselines on learned style. Given a fixed inference path, our method pareto dominates baselines for image generation both on the same category as training (left) and when evaluated on categories different from training, e.g., trained on human portraits but tested on dog images (right). We further evaluate the diversity of generated images in the supplement. We show that baselines often lose diversity, while our method leads to diverse generations while still achieving lower perceptual distance to the ground truth style. Increased marker size corresponds to an increase in the guidance scale.
Datasets.   To enable large-scale quantitative evaluation, we construct a diverse set of paired style and content images as follows. First, we generate 40 content images for each class: headshots, animals, and landscapes. When generating images in the headshot class, we generate 20 images with the prompt “A professional headshot of a man” and 20 images with the prompt
“A professional headshot of a woman”. Similarly, we split the animal class into photos of dogs and cats. To curate synthetic pairs, we then apply image editing or image-to-image translation methods to all the content images to obtain the stylized version. For each unique prompt, we choose a single paired instance as training data and hold out the other pairs with the same prompt as a test set (Same Category). For each prompt, we also choose 5 pairs from each of the other prompts as a secondary test set (Different Category). We show all our synthetic training image pairs in Appendix D. By leveraging synthetic pairs for evaluation, we can train on a single synthetic pair and test our results against held out synthetic style images. Secondly, we qualitatively compare against single artist pairs in Figure 5. Next, we describe the specific methods to create the paired dataset. First, we consider the diffusion based image editing technique LEDITS++[7] to translate images into paintings. Next, we consider Cartoonization [98], a GAN-based translation technique that aplies a cartoon-like effect. We also consider Stylized Neural Painting [106], which turns photos into painting strokes using a rendering based approach. Finally, we consider the image filtering technique posterization. We provide a more detailed description of each method for creating synthetic pairs in Appendix D.
@@ 4.2 Baselines and Evaluation Metrics
Baselines.   We compare our method against – (1) DreamBooth LoRA [36, 78] (DB LoRA), (2) Concept Sliders [23] (3) IP-adapters [100], (4) IP-adapters w/ T2I, and (5) StyleDrop [86]. DB LoRA uses only the style image and fine-tunes low-rank adapters in all the linear layers in the attention blocks of the diffusion model. We evaluate different amounts of style applications for DB LoRA using the standard LoRA scale [78] . Concept sliders presents a paired image model customization method that trains a single low-rank adapter jointly on both images, with different reconstruction losses for the style and content images. We also evaluate using the standard LoRA scale . IP-adapters is an encoder-based method that does not require training for every style and takes a style image as an extra condition separate from the text prompt. Increasing or decreasing the guidance from the input style image is possible by scaling the weight of the image conditioning. We consider the SDXL [68] implementation of this method. For the IP-Adapter, we compare against the stronger baseline of providing extra conditioning of an edge map of the content image through T2I Adapters [60] to preserve the content image structure. The recently proposed Styledrop [86] technique for learning new styles is based on MUSE [11], and uses human feedback in its method. Since MUSE is not publicly available, we follow Style-Aligned Image Generation’s [31] setup, and implement a version of StyleDrop on SDXL. Specifically, we train low-rank linear layers following each Feed-Forward layer in the attention blocks of SDXL. For a fair comparison, we train Styledrop without human feedback.
@@
Evaluation metrics.   When evaluating the performance of each method, we consider two quantitative metrics: perceptual distance to ground truth style images and structure preservation from the original image. A better customization method will have a low perceptual distance to the ground truth style images while still preserving content of the original image before adding style. We measure these using – (1) Distance to GT Styled: given holdout ground truth style images, we measure the perceptual distance between our styled outputs and the ground truth style images using DreamSim [18], a recent method for measuring the perceptual distance between images. DreamSim image embeddings are comprised of an ensemble of image embedding models, including CLIP [69] and DINO [10], which are then fine-tuned so the final embeddings respect human perception. We measure DreamSim distance as (1 - cosine similarity between DreamSim embeddings), where a lower value implies that the images are perceptually more similar. (2) Distance to Content Image: to measure content preservation after style application, we measure the perceptual distance of our generated style image to the original content image with no style guidance. We again use DreamSim, this time comparing styled and content images. Note here that a perceptual distance of zero to the content image is undesirable, as this would require no style to be applied. However, a better-performing method should obtain a better tradeoff between the two distances. (3) We also perform a human preference study of our method against baselines.